Three times Geoof's age plus 3 is Myrka's age. Myrka's is 48. What is Geoffs age?

Answer by Cromlix(4381)

Geoff's age = 3x + 3 = Myrka's age
3x + 3 = 48
3x = 48 - 3
3x = 45
x = 15
Therefore Geoff's age is 15 yrs old
Hope this helps.
:-)